# A Market for Accuracy: Classification under Competition

### v1

@@ Appendix A Proofs
Let $(x,y)\subseteq[h,b]$ be the largest consecutive interval that includes $h_{1}$ such that $\forall h^{\prime}\in(x,y)\to g(h^{\prime})<2$. Since $f_{0},f_{1}$ are continuous, then we know $g$ is continuous. Therefore, either $g(y)=2$ and $g^{\prime}(x)>0$, or $y=b$. From 5, we receive that $U_{[a,h]}(y|h)>U_{[a,h]}(h_{1}|h)$, and therefore $h_{1}$ cannot be a best response.
Proposition 3 (Convergence after 1 round).
@@
From the proof of Theorem 3, we know that $h_{i}^{1}=\underset{h\in\{a,h_{opt},P_{+}^{-1}(1/2)\}}{\operatorname*{argmax}}U(h|h_{opt})$.
@@
Assume for the sake of contradiction that $h_{j}^{1}<h_{opt}$
@@
Proposition 4 (“I improve, you improve”).
@@
From the proof of convergence in Proposition 3, we receive that in all threshold games, the players go to either side of an optimal classifier $h_{opt}$.
@@
For player $i$, from Proposition 1, $\mu_{i}=\frac{1}{2}(a_{i}+\delta_{ij})$.
$a_{i}$ remains the same, but $\delta_{ij}$ increase because $h_{j}^{1}$ went further away to the other side of $h_{opt}$, and as explained in the proof of Proposition 3, $\delta_{h_{i}^{1},opt}\subset\delta_{h_{i}^{1},h_{j}^{1}}$.
@@
Proposition 5 (Market share increases during competition).
@@
Corollary 2 (Welfare increases during competition).
This is immediate from Proposition 5 since $SW=\sum_{i}\mu_{i}$.
Proposition 6
@@
From Proposition 1 we receive that $\mu(h_{2}|h_{1})=\frac{1}{2}(a_{2}+\delta_{21})$.
@@ Appendix B Additional theoretical results
In Section 3, we mentioned that our problem setting is proven to have a PNE, a result shown by (Ben-Porat & Tennenholtz, 2019) through the use of an exact potential function. Additionaly, (Monderer & Shapley, 1996) show that every potential game is isomorphic to some congestion game; this connection however is not always readily evident. We show here the exact reduction of our problem setting to a congestion game, and highlight that the cost function is negative, which may be counterintuitive to more classic settings of congestion games.
@@
Let $s_{i}^{k}$ be the number of samples that player $i$ is accurate on along with $k-1$ other players.
@@ D.1 Main results for additional settings
Table 3 shows the learning performance of the service providers when using XGBoost as the model class for training and inference. The details of the Xgboost implementation and hyperparameters can be found in Section C.2. A comparison of interest is the general improvements of the players relative to the Linear model class. We can notice that the welfare improvement, which is equal to the total market share of all players, is significantly lower for both the Adult and COMPAS-ARREST dataset. This does not indicate that the total welfare is lower, but rather that due to the increased expressiveness of the XGboost model, the starting welfare began at a higher value.
@@
In a similar manner, Table 4 presents the results on competition where the providers are employing a Random Forest model, whose details can be found in Section C.2. We can observe that the results resemble those of the XGboost model class, which can be expected due to the similarity in nature of all Decision Tree models.
@@
As mentioned in Appendix C, each experiment was run over 10 train-test splits and the metric values were averaged out over those splits. Table 5 portrays, for each metric, the maximum variation of the standard error among the three model classes analyzed. We note that all of the errors are below 5%, and most of the errors are well below 3%.
@@ D.2 Competition Dynamics
In order to provide a more comprehensive analysis of how the market shares of the providers move through the best responses in the competition, Figure 5 shows, for each number of players competing for market share, the shifts in market share of each provider based on their positions in the game (order of play).
@@
    Order of play. In line with what is expressed in Section 6.2, The order of play when offering a best-response is important, and varies as a function of the number of players. For $n=2$ providers, playing second can offer a significant competitive advantage. When shifting to markets with $n\geq 3$ providers, however, we observe a significant advantage to the 1st mover, as discussed in Section 6.2.
@@
    General market share trend. Regardless of the order of play, and as alluded to in Table 3, the market share rises for all players, which supports the empirical claim that providers that are competing are in a way collaborating to figure out how optimally divide the market.
@@ D.3 Welfare
Figure 6 shows the social welfares across the model classes described in Section C.2, namely LinearSVC, XGBoost, and RandomForest. The test welfares are shown for the COMPAS-arrest dataset, and are portrayed for each model class and each game varying the nubmer of players. We can see that, as in Section 6.2, the welfare gets mazimized very early for the Linear model class, but hits a non-maximal plateau for the Decision Trees. This is another example of the non-monotonic nature of social welfare: in many cases providers having less expressiveness in their models will in fact benefit the consumers.
@@
The social welfare trends across different levels of data representations can also behave in a surprising manner, as shown in Section 6.2 and explained in great detail by (Jagadeesan et al., 2024). As a portrayal of this phenomenon across different competition settings, Figure 7 plots the social welfares across experiments of 2,3, and 4 players, respectively. We can observe, that while no as blatant as with $n=2$ providers, the general trend across all player formats is that the social welfare increases as an inverse proportion to the richness of the data representations.
@@ D.4 Synthetic Data
In Section 6.1, we showed an example of the chicken dynamic between asymmetrical class-conditioned gaussians. Additionally, we performed an overlap analysis between symmetric gaussians, where for each measure of distance between the means, or overlap, we calculated the best-response thresholds and performance metrics. Figure 8 (Left) shows the above analysis on asymmetric gaussians, namely $\sigma_{-1}=2,\sigma_{+1}=1$. As in Section 6.1, at each overlap the thresholds of the players are initialized at $h_{1}^{0}=h_{2}^{0}=h_{opt}$, where $h_{opt}$ is the naively optimal classifier that maximizes accuracy on the distributions. Here too, and as is guaranteed by Theorem 3, the best-response dynamics converge after just one round. In this case of asymmetry between the class-conditioned gaussians, we notice a distinctly different behavior. while $h_{1}$ remains in the same proximity to $h_{opt}$ as in the symmetric case, threshold $h_{2}$ has a “tipping point”, where it suddenly jumps to the far end of both distributions. From the accuracy graph we can also see a sudden drop in accuracy coming fromm model $h_{2}$. In regards to the market share, however, the provider that gets a spike in market share is in fact the one who played $h_{1}$ and remains close to the optimal, while the market share of $h_{2}$ simply shows a gradual increase, with no reference to a tipping point.
This phenomenon is understandable when we look at Theorem 3, that states that the best-response may be at the far end of the distributions. This occurs when either of the values $g^{-1}(1/2),g^{-}1(2)$ stops existing, where $g(x)=\frac{f_{1}(x)}{f_{0}(x)}$ is the MLR function. In these cases, the left or right threshold (depending on which value of $g^{-1}$ disappears) will continue to gain by moving to the far end of the interval. This is precisely what is shown in Figure 8 (Right); at a certain overlap ($\sim-1$), there is no threshold $h$ for which $g(h)=2$, and so the gain in discrepancy for $h_{2}$ will continue to outweigh the loss in accuracy, and $h_{2}$ ends up at the far right end. The market share of $h_{1}$ is then suddenly increased, since while its accuracy remains the same, the discrepancy from $h_{2}$ is now significantly greater.
@@ D.5 Asymmetrical power between players
Another interesting question we can ask from the perspective of the learners/providers is, if a provider were to invest cost and effort to gain better data, how would this help them in competition? In naïve settings, i.e. single-provider markets, the answer to this is straightforward: more data = better accuracy. In accuracy markets, however, the specialization needs to be considered as well, and as we have seen (Section 6.2), the behavior of the market when altering the data quality can be counter-intuitive. To measure the gain to be had when improving data, we ran experiments where one of the providers has access to more features than its counterpart/s. For each of 2 possible move positions (1st or 2nd), two metrics were considered:
@@
Table 6 shows the above metrics for markets with 2,3, and 4 providers, and for various possibilities of data improvement.
@@
    $\Delta\mu$ from regular setting. One of our central insights from Section 6.2 is that when $n=2$ providers, moving second is beneficial, and when $n>2$, the opposite is true, and moving first is better. In the case of investing in better data, and when comparing the provider’s market gain vs. himself in a regular setting, we see an inverse effect.
    For $n=2$, better data creates market gains only when you are the first mover, as in certain lopsided markets the gain is $>12\%$. For example, in the case where the better-data provider has 15 features, and the other providers have 3 features, we can observe a 12% gain, which measures the benefit of investing in the additional 12 features.
@@
    $\Delta\mu$ from the next-best provider. When comparing the difference across data-variation experiments in market shares between providers, we notice that the trend behaves similarly to how we have seen in the order-of-play results of Section 6.2. We can observe that, interestingly, if a provider improved his absolute market-share relative to himself, this doesn’t translate to the provider improving his market share relative to others. Take for example the cases where $n>2$ and the provider moves 2nd. As stated above and as can be seen in Table 6, the added data advantage in this setting helps the provider gain in absolute market share. The market gain relative to the other providers, however, has an inverse result, and in many cases the competitors end up with a better overall utility. This tells us that in these settings, when the advantaged provider goes second, the total welfare (as the sum of individual market shares) increases.

### v2

@@ Appendix A Proofs
Let $(x,y)\subseteq[h,b]$ be the largest consecutive interval that includes $h_{1}$ such that $\forall h^{\prime}\in(x,y)\to g(h^{\prime})<2$. Since $f_{0},f_{1}$ are continuous, then we know $g$ is continuous. Therefore, either $g(y)=2$ and $g^{\prime}(x)>0$, or $y=b$. From ˜5, we receive that $U_{[a,h]}(y|h)>U_{[a,h]}(h_{1}|h)$, and therefore $h_{1}$ cannot be a best response.
Proposition˜3 (Convergence after 1 round).
@@
From the proof of Theorem˜3, we know that $h_{i}^{1}=\underset{h\in\{a,h_{opt},P_{+}^{-1}(1/2)\}}{\operatorname*{argmax}}U(h|h_{opt})$.
@@
Assume for the sake of contradiction that $h_{j}^{1}<h_{opt}$ -
@@
Proposition˜4 (“I improve, you improve”).
@@
From the proof of convergence in Proposition˜3, we receive that in all threshold games, the players go to either side of an optimal classifier $h_{opt}$.
@@
For player $i$, from Proposition˜1, $\mu_{i}=\frac{1}{2}(a_{i}+\delta_{ij})$.
$a_{i}$ remains the same, but $\delta_{ij}$ increase because $h_{j}^{1}$ went further away to the other side of $h_{opt}$, and as explained in the proof of Proposition˜3, $\delta_{h_{i}^{1},opt}\subset\delta_{h_{i}^{1},h_{j}^{1}}$.
@@
Proposition˜5 (Market share increases during competition).
@@
Corollary˜2 (Welfare increases during competition).
This is immediate from Proposition˜5 since $SW=\sum_{i}\mu_{i}$.
Proposition˜6
@@
From Proposition˜1 we receive that $\mu(h_{2}|h_{1})=\frac{1}{2}(a_{2}+\delta_{21})$.
@@ Appendix B Additional theoretical results
In Section˜3, we mentioned that our problem setting is proven to have a PNE, a result shown by (Ben-Porat & Tennenholtz, 2019) through the use of an exact potential function. Additionaly, (Monderer & Shapley, 1996) show that every potential game is isomorphic to some congestion game; this connection however is not always readily evident. We show here the exact reduction of our problem setting to a congestion game, and highlight that the cost function is negative, which may be counterintuitive to more classic settings of congestion games.
@@
Let $s_{i}^{k}$ be the number of samples that player $i$ is accurate on along with k1k-1 other players.
@@ D.1 Main results for additional settings
Table˜3 shows the learning performance of the service providers when using XGBoost as the model class for training and inference. The details of the Xgboost implementation and hyperparameters can be found in Section˜C.2. A comparison of interest is the general improvements of the players relative to the Linear model class. We can notice that the welfare improvement, which is equal to the total market share of all players, is significantly lower for both the Adult and COMPAS-ARREST dataset. This does not indicate that the total welfare is lower, but rather that due to the increased expressiveness of the XGboost model, the starting welfare began at a higher value.
@@
In a similar manner, Table˜4 presents the results on competition where the providers are employing a Random Forest model, whose details can be found in Section˜C.2. We can observe that the results resemble those of the XGboost model class, which can be expected due to the similarity in nature of all Decision Tree models.
@@
As mentioned in Appendix˜C, each experiment was run over 10 train-test splits and the metric values were averaged out over those splits. Table˜5 portrays, for each metric, the maximum variation of the standard error among the three model classes analyzed. We note that all of the errors are below 5%, and most of the errors are well below 3%.
@@ D.2 Competition Dynamics
In order to provide a more comprehensive analysis of how the market shares of the providers move through the best responses in the competition, Figure˜5 shows, for each number of players competing for market share, the shifts in market share of each provider based on their positions in the game (order of play).
@@
    Order of play. In line with what is expressed in Section˜6.2, The order of play when offering a best-response is important, and varies as a function of the number of players. For $n=2$ providers, playing second can offer a significant competitive advantage. When shifting to markets with $n\geq 3$ providers, however, we observe a significant advantage to the 1st mover, as discussed in Section˜6.2.
@@
    General market share trend. Regardless of the order of play, and as alluded to in Table˜3, the market share rises for all players, which supports the empirical claim that providers that are competing are in a way collaborating to figure out how optimally divide the market.
@@ D.3 Welfare
Figure˜6 shows the social welfares across the model classes described in Section˜C.2, namely LinearSVC, XGBoost, and RandomForest. The test welfares are shown for the COMPAS-arrest dataset, and are portrayed for each model class and each game varying the nubmer of players. We can see that, as in Section˜6.2, the welfare gets mazimized very early for the Linear model class, but hits a non-maximal plateau for the Decision Trees. This is another example of the non-monotonic nature of social welfare: in many cases providers having less expressiveness in their models will in fact benefit the consumers.
@@
The social welfare trends across different levels of data representations can also behave in a surprising manner, as shown in Section˜6.2 and explained in great detail by (Jagadeesan et al., 2024). As a portrayal of this phenomenon across different competition settings, Figure˜7 plots the social welfares across experiments of 2,3, and 4 players, respectively. We can observe, that while no as blatant as with $n=2$ providers, the general trend across all player formats is that the social welfare increases as an inverse proportion to the richness of the data representations.
@@ D.4 Synthetic Data
In Section˜6.1, we showed an example of the chicken dynamic between asymmetrical class-conditioned gaussians. Additionally, we performed an overlap analysis between symmetric gaussians, where for each measure of distance between the means, or overlap, we calculated the best-response thresholds and performance metrics. Figure˜8 (Left) shows the above analysis on asymmetric gaussians, namely $\sigma_{-1}=2,\sigma_{+1}=1$. As in Section˜6.1, at each overlap the thresholds of the players are initialized at $h_{1}^{0}=h_{2}^{0}=h_{opt}$, where $h_{opt}$ is the naively optimal classifier that maximizes accuracy on the distributions. Here too, and as is guaranteed by Theorem˜3, the best-response dynamics converge after just one round. In this case of asymmetry between the class-conditioned gaussians, we notice a distinctly different behavior. while $h_{1}$ remains in the same proximity to $h_{opt}$ as in the symmetric case, threshold $h_{2}$ has a “tipping point”, where it suddenly jumps to the far end of both distributions. From the accuracy graph we can also see a sudden drop in accuracy coming fromm model $h_{2}$. In regards to the market share, however, the provider that gets a spike in market share is in fact the one who played $h_{1}$ and remains close to the optimal, while the market share of $h_{2}$ simply shows a gradual increase, with no reference to a tipping point.
This phenomenon is understandable when we look at Theorem˜3, that states that the best-response may be at the far end of the distributions. This occurs when either of the values $g^{-1}(1/2),g^{-}1(2)$ stops existing, where $g(x)=\frac{f_{1}(x)}{f_{0}(x)}$ is the MLR function. In these cases, the left or right threshold (depending on which value of $g^{-1}$ disappears) will continue to gain by moving to the far end of the interval. This is precisely what is shown in Figure˜8 (Right); at a certain overlap ($\sim-1$), there is no threshold $h$ for which $g(h)=2$, and so the gain in discrepancy for $h_{2}$ will continue to outweigh the loss in accuracy, and $h_{2}$ ends up at the far right end. The market share of $h_{1}$ is then suddenly increased, since while its accuracy remains the same, the discrepancy from $h_{2}$ is now significantly greater.
@@ D.5 Asymmetrical power between players
Another interesting question we can ask from the perspective of the learners/providers is, if a provider were to invest cost and effort to gain better data, how would this help them in competition? In naïve settings, i.e. single-provider markets, the answer to this is straightforward: more data = better accuracy. In accuracy markets, however, the specialization needs to be considered as well, and as we have seen (Section˜6.2), the behavior of the market when altering the data quality can be counter-intuitive. To measure the gain to be had when improving data, we ran experiments where one of the providers has access to more features than its counterpart/s. For each of 2 possible move positions (1st or 2nd), two metrics were considered:
@@
Table˜6 shows the above metrics for markets with 2,3, and 4 providers, and for various possibilities of data improvement.
@@
    $\Delta\mu$ from regular setting. One of our central insights from Section˜6.2 is that when $n=2$ providers, moving second is beneficial, and when $n>2$, the opposite is true, and moving first is better. In the case of investing in better data, and when comparing the provider’s market gain vs. himself in a regular setting, we see an inverse effect.
    For $n=2$, better data creates market gains only when you are the first mover, as in certain lopsided markets the gain is $12\%$. For example, in the case where the better-data provider has 15 features, and the other providers have 3 features, we can observe a 12% gain, which measures the benefit of investing in the additional 12 features.
@@
    $\Delta\mu$ from the next-best provider. When comparing the difference across data-variation experiments in market shares between providers, we notice that the trend behaves similarly to how we have seen in the order-of-play results of Section˜6.2. We can observe that, interestingly, if a provider improved his absolute market-share relative to himself, this doesn’t translate to the provider improving his market share relative to others. Take for example the cases where $n>2$ and the provider moves 2nd. As stated above and as can be seen in Table˜6, the added data advantage in this setting helps the provider gain in absolute market share. The market gain relative to the other providers, however, has an inverse result, and in many cases the competitors end up with a better overall utility. This tells us that in these settings, when the advantaged provider goes second, the total welfare (as the sum of individual market shares) increases.